Pump Up Gym has an initial joining fee of $205 and monthly membership dues of $15. Universe Gym has an initial joining fee of $125 and monthly membership dues of $19.
18.
When will the costs to join and maintain membership at the gyms be equal?

• At 20 months
• At 15 months
• At 22 months
• At 18 months





The point (0, 2, –6) lies in the yz-plane and is located to the right and down from the origin. What is a similar description for each point?
23.
(3, –5, 0)
Lies on the xy-plane and is located backward and to the left of the origin.
Lies on the xy-plane and is located forward and to the left of the origin.
Lies on the xy-plane and is located forward and to the right of the origin.
Lies on the xy-plane and is located backward and to the right of the origin.





Answer by Alan3354(21610)

Pump Up Gym has an initial joining fee of $205 and monthly membership dues of $15. Universe Gym has an initial joining fee of $125 and monthly membership dues of $19.
18.
When will the costs to join and maintain membership at the gyms be equal?
p = 205+15m
u = 125+19m
At p = u:
205+15m = 125+19m
80 = 4m
m = 20 months
------------------------------
The point (0, 2, –6) lies in the yz-plane and is located to the right and down from the origin. What is a similar description for each point?
23.
(3, –5, 0)
Lies on the xy-plane and is located forward and to the left of the origin.